OpenStudy (anonymous):

A weather satellite in circular orbit around Earth completes one orbit every 5 hours. The radius of Earth is about 6,400 km and the satellite is positioned 4,700 km above the Earth. How far does the satellite travel in 1 hour? Round your answer to the nearest kilometer.

OpenStudy (anonymous):

@algebra2sucks :you there ?

OpenStudy (anonymous):

Yes lol sorry

OpenStudy (anonymous):

np ,do you know what is angular velocity ?

OpenStudy (anonymous):

nope :/

OpenStudy (anonymous):

Hmmm ,First of all you should know that the angular velocity is the same . \[\omega= {2\pi\over T}={2\pi\over 5}{\rm rad/hr}\\ v=r\times\omega=?\] Then , the distance travelled in 1hr =\[v\times 1hr\]

OpenStudy (anonymous):

Can you continue now ?

OpenStudy (anonymous):

do we use 6,400 and 4,700

OpenStudy (anonymous):

Yes ,of course ==> (r)

OpenStudy (anonymous):

whats the w? so 6,400 x what?

OpenStudy (anonymous):

\[\Large \color{Red}{\bf \omega=angular~velocity~.}\] \[\Large v=r\times\omega\] \[\Large \color{MidnightBlue}{\omega={2\pi\over 5}{\rm rad/hr}\\ v=(6400+4700){2\pi\over 5}=11100\times {2\pi\over 5}{\rm km/hr}\\ d=v\times t=11100\times{2\pi\over 5}{\rm km/hr}\times1{\rm hr}\\ \Large\boxed{d=??~{\rm km}}}\] Calculate the d (Distance) . ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ Got it ?

OpenStudy (anonymous):

Yes i got it!

OpenStudy (anonymous):

Well i little i got it lol

OpenStudy (anonymous):

lol ,If you have any questions ,Feel free to ask :)

OpenStudy (anonymous):

do I calculate 2 divided by 5 then times the answer?

OpenStudy (anonymous):

\[\Large \bf d=11100 \times \frac{2 \times \pi}{5} \times 1 \]

OpenStudy (anonymous):

5,906 or 8,042? for the answer?

OpenStudy (anonymous):

@Eyad